OpenStudy (anonymous):

A video shop owner wants to find out how many videos are turned in late. He found that of the next 35 videos turned in, 7 were late. If 560 videos are turned in over the next several hours, what is the best prediction for how many will be late? A. 110 B. 115 C. 113 D. 112

OpenStudy (lstrasmann):

\[\frac{ 7 }{ 35 }=\frac{ x }{ 560 }\]

OpenStudy (campbell_st):

use ratios late/total 7/35 = x/560

OpenStudy (lstrasmann):

Solve for x ^

OpenStudy (anonymous):

Can you help im not sure how to do that sorry

OpenStudy (campbell_st):

well 7/35 = 1/5 so find 1/5 of 560...

OpenStudy (lstrasmann):

Okay, this is a proportion. You cross-multiply to find the answer.